OpenStudy (anonymous):

the x-intercept of a line is -5 and the y-intercept of the line is -2. what is the equation of the line? A. y= -5/2x-5 B. y= 2/5x+2 C. y= 5/2x-2 D. y= -2/5x-2

ganeshie8 (ganeshie8):

use the intercept form of line :- \(\large \frac{x}{a} + \frac{y}{b} = 1\)

ganeshie8 (ganeshie8):

and simplify it to y = mx+b form

ganeshie8 (ganeshie8):

in above equation/s, a = x intercept b = y intercept

OpenStudy (anonymous):

I use to know how to do this but im so confused now

ganeshie8 (ganeshie8):

its okay, let me show u how to simplify

OpenStudy (anonymous):

alright thank you

ganeshie8 (ganeshie8):

\(\large \frac{x}{-5} + \frac{y}{-2} = 1\) add x/5 both sides \(\large \frac{y}{-2} = \frac{x}{5} + 1\) multiply -2 thru out \(\large y = \frac{-2x}{5} - 2\)

ganeshie8 (ganeshie8):

see if that makes some sense

OpenStudy (anonymous):

so your adding x/5 to the other side oooh I see now because multiplying -2 to them

OpenStudy (anonymous):

that made so much sense thank you

ganeshie8 (ganeshie8):

glad to hear :))

OpenStudy (anonymous):

:) thank you :D

ganeshie8 (ganeshie8):

np :)